OpenStudy (anonymous):

Simplify each quotient. sqrt{3}(sqrt{12}+4)

OpenStudy (anonymous):

sqrt(12) = 2*sqrt(3) sqrt(3)*sqrt(3)*2 = 6 6/4 = 3/2 = 1.5

OpenStudy (anonymous):

\[\sqrt{3}(\sqrt{12}+4)\] \[\sqrt{3}\sqrt{12}+\sqrt{3}*4\] \[3^{1/2}*12^{1/2}+\sqrt{3}*4\] \[(3*12)^{1/2}+\sqrt{3}*4\] \[36^{1/2}+\sqrt{3}*4 = \sqrt{36}+\sqrt{3}*4\]

OpenStudy (anonymous):

\[\sqrt{36}=6\] \[6+\sqrt{3}*4\]